The volume of a refrigerator is x^3-3x^2-16x-12. Its height is x + 2. What are the other two dimensions?
Answer by ewatrrr(24785)

 
Hi,
Note: V(rectangular solid) = l*w*h
The volume of a refrigerator is x^3-3x^2-16x-12, height = (x+2)
(x^3-3x^2-16x-12)/(x+2)= x^2 - 5x - 6 = (x-6)(x+1)
Using Synthetic Division:
-2 1 -3 -16 -12
2 10 12
1 -5 -6 0
V = (x+2)(x-6)x+1)
The other two dimensions are: (x-6)and (x+1)